# Conformal prediction of future insurance claims in the regression problem

### v2

@@ 5.2 Personal injury insurance claims data
total settled amount \$10\$4,490,000currency-dollar10currency-dollar4490000\\$10–\\$4,490,000\$ 10 – \$ 4 , 490 , 000
@@
op__\__time operational time 0.1–99.1
@@
Except for the variable “total”, all other variables are categorical and have been assigned numerical scores. For variables “inj1”, “inj2”, “inj3”, “inj4”, and “inj5”, there are lots of missing values. We assign 00 to each of these missing values. Table 6 displays summary statistics of all these variables.
@@
We apply the proposed method to this data set with ‘total’ being the response variable and all other variables being predictors. Table 7 shows the upper bound of the $100(1-\alpha)\%$ conformal prediction interval in Theorem 2 for three $\alpha$ levels: $0.90,0.95$, and $0.995$. Given the excellent performance of the proposed conformal prediction interval in Examples 1–3, we are inclined to believe that the conformal prediction interval here achieves the nominal coverage probability level for each $\alpha$ value. In particular, if the insurer wants to comply with Solvency II, they can set their risk capital level to $544,093$ for this line of business.
@@ 6 Concluding remarks
Explanation and prediction are the two major tasks in data science and statistics. When it comes to actuarial science, prediction is usually the main goal. A finite-sample valid prediction interval is crucial for an insurer to understand the risk associated with their business and to set risk capital at an appropriate level. In particular, the insurer can apply it to comply with the risk capital requirements in Solvency II, which dictates that the insurer must be able to cover losses with a $99.5\%$. This article proposed a new non-conformity measure to create a new prediction interval for the response variable in the regression problem. This prediction interval has multiple merits: (i) it avoids both model misspecification risk and the selection effect, (ii) it has a closed-form formula and hence is easy to compute, and (iii) it guarantees finite-sample validity. Our simulation study shows that this prediction interval achieves finite-sample validity at the confidence level $99.5\%$ across different scenarios. In this paper, we have restricted the domain of the response variable $Y$ to $\mathbb{R}_{+}$ since the response variable is non-negative in most insurance applications. However, it is easy to see that the results in this paper extend to the case where $Y\in\mathbb{R}$. Finally, the method proposed in this paper does not apply to the case of a categorical response variable. To apply comformal prediction to predict a categorical response variable would require a different strategy than the one proposed in this article.

### v3

@@ 5.2 Personal injury insurance claims data
total settled amount $\mathdollar 10–\mathdollar 4,490,000$
@@
op_\_time operational time 0.1–99.1
@@
Except for the variable “total”, all other variables are categorical and have been assigned numerical scores. For variables “inj1”, “inj2”, “inj3”, “inj4”, and “inj5”, there are lots of missing values. We assign 0 to each of these missing values. Table 6 displays summary statistics of all these variables.
@@
We apply the proposed method to this data set with ‘total’ being the response variable and all other variables being predictors. Table 7 shows the upper bound of the $100(1-\alpha)\%$ conformal prediction interval in Theorem 2 for three $\alpha$ levels: $0.90,0.95$, and $0.995$. Although we do not know the true distribution here, we know from Section 3 that a conformal prediction interval is provably valid; hence the three conformal prediction intervals in Table 6 all achieve the nominal coverage probability level for their respective $\alpha$ value. In particular, if the insurer wants to comply with Solvency II, they can set their risk capital level to $544,093$ for this line of business.
@@ 6 Concluding remarks
Explanation and prediction are the two major tasks in data science and statistics. When it comes to actuarial science, prediction is usually the main goal. As Example 2 demonstrates, if a prediction interval is created using a parametric model (e.g., a generalized linear model), it can undercover when the proposed model fails to match the true data-generating mechanism. This model misspecification issue is avoided when a prediction interval is based on a non-parametric method (e.g., Random Forests). However, such a prediction interval may not be finite-sample valid. For example, Examples 1 to 3 show that Random Forests prediction intervals are not finite-sample valid. In practice, a finite-sample valid prediction interval is crucial for an insurer to understand the risk associated with their business and to set risk capital at an appropriate level. In particular, the insurer can apply it to comply with the risk capital requirements in Solvency II, which dictates that the insurer must be able to cover losses with a $99.5\%$. This article proposed a new non-conformity measure to create a new prediction interval for the response variable in the regression problem. This prediction interval has multiple merits: (i) it avoids both model misspecification risk and the selection effect, (ii) it has a closed-form formula and hence is easy to compute, and (iii) it guarantees finite-sample validity. Our simulation study shows that this prediction interval achieves finite-sample validity at the confidence level $99.5\%$ across different scenarios. In this paper, we have restricted the domain of the response variable $Y$ to $\mathbb{R}_{+}$ since the response variable is non-negative in most insurance applications. However, it is easy to see that the results in this paper extend to the case where $Y\in\mathbb{R}$. Finally, the method proposed in this paper does not apply to the case of a categorical response variable. To apply conformal prediction to predict a categorical response variable would require a different strategy than the one proposed in this article.
Acknowledgments
I thank the Editor for his support and an anonymous reviewer for many helpful comments and suggestions.